OpenStudy (anonymous):

Use a table of values to find the solution to the equation ƒ(x ) = h(x ) where f(x)=1/1x and h(x ) = 2x + 3. Show your work

OpenStudy (anonymous):

you know the values of functions and so you need to solve the equation 1/x=2x+3

OpenStudy (anonymous):

is 1/1x the 1/x?

OpenStudy (anonymous):

1/1x and 1/x is he same @ksanka

OpenStudy (anonymous):

@ksanka can you please help me?

OpenStudy (perl):

1/x

OpenStudy (anonymous):

isolate the variable. 1/x-2x-3=0 then \[(1-2x^2-3x)/x=0\]

OpenStudy (anonymous):

-2x^2-3x+1=0 \[2x^2+3x-1=0\] use quadratics formula to find the roots

OpenStudy (perl):

1/x=2x+3 1= 2x^2 + 3x 2x^2 + 3x - 1 = 0

OpenStudy (anonymous):

\[x=\frac{ b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

use this formula to calculate the roots.

OpenStudy (anonymous):

there has to be two solutions. 

OpenStudy (perl):

it says to use a table of values, do you mean

OpenStudy (anonymous):

i really dont understand what that means either @perl

OpenStudy (perl):

you could graph a table of values and see where they are closest

OpenStudy (perl):

for what x value are the y values closest

OpenStudy (anonymous):

Im not sure how to do that @perl

OpenStudy (anonymous):

I think they mean to create a table of values and find the value of x when f(x)=h(x)

OpenStudy (anonymous):

so first calclulate the roots and then create a table with the values of x close to the roots